OpenStudy (anonymous):

HELP ME bumped this question too many times CDE~QRS and has a ratio of similarity of 4:1. If side RS measures 8 inches, how long is side DE? IF ITS WRONG JUST GIVE ME THe correct formula my answer 2

OpenStudy (lgbasallote):

this is a triangle right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

and which one is longer? CDE? or QRS?

OpenStudy (anonymous):

well theres no picture but im guessing its QRS

OpenStudy (anonymous):

32 and CDE is larger

OpenStudy (lgbasallote):

if there's no pictirue then CDE:QRS correspondss to the ratio 4:1 which is bigger...4 or 1?

OpenStudy (anonymous):

naturally 4 i dont get were your gettin g at

OpenStudy (anonymous):

but i really havent gotten the rules straight with these problems

OpenStudy (lgbasallote):

and 4 is on the left side of 4:1 right? on CDE:QRS which is on the left side?

OpenStudy (lgbasallote):

you'll see where im getting at later on ;D

OpenStudy (anonymous):

DE:RS = 4:1 DE/8 = 4/1 DE = 8*4 DE=32

OpenStudy (anonymous):

the ratio means that lengths of sides of triangle CDE is 4 times that of PQR

OpenStudy (anonymous):

i did crooss multiply thinking id get the right answer (2) guess not thanks